OpenStudy (anonymous):

which number is greater ? 111111111 in base 2 or 4321 in base 5

OpenStudy (anonymous):

Can u convert 4321 base 5 into base 2

OpenStudy (anonymous):

That's going to help me get the answer ?

OpenStudy (anonymous):

Of course

OpenStudy (anonymous):

I don't understand how to do that though

OpenStudy (unklerhaukus):

\[5^5,5^4,5^3,5^2,5^1,5^0=3125,625,125,25,5,1\]

OpenStudy (anonymous):

4321 in base 2 is 1001001010

OpenStudy (anonymous):

Sorry, But I got 10001101110

OpenStudy (unklerhaukus):

4321 1*3125 _ = 1196 1196 1*625 _ = 571 571 4*125 _ = 71 71 2*25 _ =21 21 4*5 _ =1 1 1*1_ =0 \[4321= 114241_5 \]

OpenStudy (unklerhaukus):

numbers do not change size ,

OpenStudy (anonymous):

What was that @UnkleRhaukus

OpenStudy (unklerhaukus):

i just converted 4321 into base 5

OpenStudy (anonymous):

4321 is in base 5 we need to convert it into base 2

ganeshie8 (ganeshie8):

111111111 = 2^9-1 in decimal

OpenStudy (unklerhaukus):

bother